OpenStudy (anonymous):

Please help!! Leah is purchasing a condominium and is financing $515,000 with a 20-year 5/1 ARM at 5.25% with a 1/12 cap structure. What will her payments be at the beginning of year 6?

OpenStudy (amistre64):

my finance speak is a bit rusty; can you define the terms of the loan in laymans terms? what does the 5/1 tells us? and what is a 1/12 cap structure?

OpenStudy (anonymous):

I have no idea

OpenStudy (amistre64):

but that is part of your responsibility to know .... if we can iron out what they tell us about the loan, calculating the rest of it would be simple enough

OpenStudy (anonymous):

Well I don't know how to solve these kinda of problems that is why I am asking for help

OpenStudy (amistre64):

how to read the information is important to know in order to develop a strategy for a solution. As is, the cap structure seems off; the format tends to be a 3 value setup so you might want to check to see if that posted correctly

OpenStudy (anonymous):

It is.

OpenStudy (amistre64):

that at least tells us how to intepret the information ... how to define payments for it tho, hmmm you wouldnt happen to have a formula related to this would you? im used to seeing this stuff as a fixed payment if anything, the remianing balance of the loan is restructed for the remaining time limit at the new interest rate

OpenStudy (amistre64):

this gives me an idea at least :)

OpenStudy (anonymous):

K cool :)

OpenStudy (amistre64):

the payments on a 20 year loan at 5.25 would amount to: 515,000(1+.0525/12)^(20*12) = x(1-(1+.0525/12)^(20*12))/(-.0525/12)), which the wolf says is 3470.30 cents calculating the balance after 5 years: 515,000(1+.0525/12)^(20*12) -1370.30(1-(1+.0525/12)^(20*12))/(-.0525/12) gives us a balance of 575,422 to restructure for the remaining 15 years at a bump up of 1% interest 575,422(1+.0625/12)^(15*12) = x(1-(1+.0625/12)^(15*12))/(-.0625/12) gives me a payment of 4933.80

OpenStudy (amistre64):

essentially: determine value of payment on the 20 year loan determine the balance remaining after 5 years restructure the loan at +1% at a 15 year time frame to determine the new payment amount

OpenStudy (amistre64):

i cant say that im 100% sure that this is right ... but it feels right to me

OpenStudy (anonymous):

These are the answer choices. $3,155.38 $3,701.45 $3,470.30 $2,908.95

OpenStudy (amistre64):

i have an error in by 5 year balance

OpenStudy (amistre64):

excel helped out :) 3701.45

OpenStudy (anonymous):

Correct answer: $3701.45